# UNB StepUP: A footStep database for gait analysis and recognition using Underfoot Pressure

### v2

@@ Acknowledgements
{forest}
for tree= font=, grow’=0, folder indent=.9em, folder icons, edge= dotted [StepUP-P150 [participant_metadata.csv, is file] [001 [BF [S1 [trial.{npz,mat},is file] [preprocessed.{npz,mat},is file] ] [S2] [S3] [W1 [metadata.csv, is file] [trial.{npz,mat}, is file] [pipeline_1.{npz,mat}, is file] [pipeline_2.{npz,mat}, is file] ] [W2] [W3] [W4] ] [ST] [P1] [P2] ] [002] [003] [
⋮, is file] [150] ]
Figure 11: Data structure for the UNB StepUP-P150 dataset.
@@
Figure 15: Comparison of four spatiotemporal gait parameters between the CASIA-D barefoot database and barefoot samples from the UNB StepUP-P150 dataset. Note: Significant differences were found between the two CASIA-D and StepUP-P150 datasets in step length, cadence, and toe-out angle, for both the preferred and fast walking trials ($p<0.05$ for all, using two-sample $t$-tests). No significant differences were found for step width.
Figure 16: Average step length and step width (also called support base) for each participant while wearing two personal pairs of shoes from the (a) CASIA-D shod and (b) StepUP-P150 databases. Note: Dotted lines connect markers associated with the same participant, with the two shoe types represented by diamond and star markers, respectively.
Table 1: Key features of public pressure-based gait databases and datasets, including those featuring healthy subjects.
@@
Footwear Other Covariates
CASIA-D [9] 88a 2,640 2-4 66M 20F 20-60 Asian PS, FS BF NA
CASIA-D [9] 30b 540 2-4 24M 6F 20-40 Asian PS BF, PM NA
SFootBD [13, 10] 127 19,980 2 83M 44F 23.7 (18-32) NA (University students in Spain) PS BF, PM Time elapsed, Load carriage
CAD WALK [11] 55 2,640 1 21M 34F 42.3 (18-70) White (Dutch) PS BF NA
UoM-Gait-69 [12] 69 14,394 3-4 32M 37F 36.4 (20-63) NA PS, FS PO Dual-task gait
UNB StepUP-P150 150 250,000 4-6 72M 78F 34 (19-91) White, Asian, Others PS, FS, SS, STS BF, PM, PC see Additional Covariates section
@@
b Downloadable resources only include data from 15 participants, of whom 13 tried two different shoe types, while the remaining 2 tried only one type.
M: male, F: female, PS: preferred speed, FS: faster than preferred speed, SS: slower than preferred speed, STS: slow-to-stop, BF: barefoot, PO: personal shoes (one pair per participant), PM: personal shoes (multiple pairs per participant), PC: controlled shoes.
@@
None No spatial normalization
Zero Padding Pad border of footstep with zeros to a specified tensor width and length
Resize Apply spatial interpolation to resize the footstep to a specified tensor width and length
Foot Rotation Rotate footstep according to the direction of its first principal component axis (the sole’s longest dimension)
@@
Registration Linearly transform (translate, rotate, scale) the footstep to align with a reference template (e.g., the MUN104 healthy barefoot templates [26])
@@
None No temporal normalization
Zero Padding Pad the 3D tensor with zeros at the end of the footstep to a specified number of frames
Interpolation Interpolate the 3D tensor to a specified number of frames (e.g., 101 frames, where each frame represents 1% of the stance)
@@
None No amplitude normalization
Body Mass Linearly rescale amplitudes by the participant’s measured body mass, so that amplitudes are relative to body weight
Mean GRF Linearly rescale amplitudes by the average ground reaction force of the footstep
Min-Max Linearly rescale amplitudes to a maximum value of 1 and minimum of 0
@@
ParticipantID 001-150 Unique identifier.
@@
Gender Man, Woman, Non-Binary, Unknown Gender identity.
@@
NeurologicalCondition Yes, No Experiencing a neurological condition (e.g., Parkinson’s, stroke).
NeurologicalConditionComment Text Comment if selected ‘Yes’ for NeurologicalCondition.
@@
Concussion Yes, No Recent concussion (<6 months).
ConcussionComment Text Comment if selected ‘Yes’ for Concussion.
Pregnancy Yes, No Pregnant at time of collection.
PregnancyComment Text Comment if selected ‘Yes’ for Pregnancy.
@@
[b] Field Possible Values/Format Description ParticipantID 001-150 Unique participant identifier Footwear BF, ST, P1, P2 Footwear condition for trial Speed W1, W2, W3, W4 Walking speed condition for trial FootstepID* 0-$N_{steps}$ Footstep’s index in 90-second trial PassID* 0-$N_{passes}$ The pass within which the footstep occurred (i.e., incremented each time the participant steps off the tiles to turn around) StartFrame* 0-$N_{frames}$ Frame in 90-second recording where footstep’s first pressure contact occurred EndFrame* 0-$N_{frames}$ Frame in 90-second recording where footstep’s pressure contact ended Ymin, Ymax* 0-719 Footstep bounding box limits along tile grid $y$-axis (parallel to walking direction, 3.6 m length) Xmin, Xmax* 0-239 Footstep bounding box limits along tile grid $x$-axis (perpendicular to walking direction, 1.2 m width) Orientation 0, 1 Footstep’s orientation on the tile grid: 1 if walking toward Tiles 11 and 12, or 0 if walking toward Tiles 1 and 2. Side Left, Right Whether the footstep corresponds to the right or left foot Standing 0, 1 1 the footstep corresponds to standing behaviour during the Slow-to-Stop (W2) trials, 0 otherwise Incomplete 0, 1 1 if the footstep was not captured fully by the sensors (e.g., partially off of the tile-grid or cut-off at beginning or end of recording), 0 otherwise Rscore 0-89 Footstep’s R-score, approximating the number of standard deviations from the trial mean Outlier 0, 1 1 if the footstep’s R-Score exceeds the recommended threshold of 2, or 0 otherwise Exclude 0, 1 The combination of the ‘Standing’, ‘Incomplete’ and ‘Outlier’ columns for easy exclusion of these footsteps if desired RotationAngle -90°- 90° Footstep’s rotation angle in degrees with respect to the tile grid’s long axis ($y$-axis) FootLength 1-75 Length of the footstep in pixels, measured along the footstep’s first principal component axis (longest foot dimension) FootWidth 1-40 Width of the footstep in pixels, measured along the footstep’s second principal component axis (perpendicular to longest foot dimension) MeanGRF Floating Point Average of footstep’s ground reaction force (total sum of all pressure values in kPa at each time point) over the duration of the stance.

### v3

@@ Acknowledgements
EM: Methodology, Investigation, Data Curation, Writing - Review & Editing.
@@
{forest}
for tree= font=, grow’=0, folder indent=.9em, folder icons, edge= dotted [StepUP-P150 [participant_metadata.csv, is file] [{py,mat} [001 [BF [S1 [trial.{npz,mat},is file] [preprocessed.{npz,mat},is file] ] [S2] [S3] [W1 [metadata.csv, is file] [trial.{npz,mat}, is file] [pipeline_1.{npz,mat}, is file] [pipeline_2.{npz,mat}, is file] ] [W2] [W3] [W4] ] [ST] [P1] [P2] ] [002] [003] [
⋮, is file] [150] ] ]
Figure 9: Data structure for the UNB StepUP-P150 dataset.
@@
Figure 15: Comparison of four spatiotemporal gait parameters between the CASIA-D barefoot database and barefoot samples from the UNB StepUP-P150 dataset. Note: Significant differences were found between the CASIA-D and StepUP-P150 datasets in step length, cadence, and toe-out angle, for both the preferred and fast walking trials ($p<0.05$ for all, using two-sample $t$-tests). No significant differences were found for step width.
Figure 16: Average step length and step width for each participant while wearing two personal pairs of shoes from the (a) CASIA-D shod and (b) StepUP-P150 databases. Note: Dotted lines connect markers associated with the same participant, with the two shoe types represented by diamond and star markers, respectively.
Table 1: Key features of public pressure-based gait databases and datasets.
@@
Footwear
CASIA-D Barefoot [9] 88a 2,640 2-4 66M 20F 20-60 Asian (Chinese) PS, FS BF
CASIA-D Shod [9] 30b 540 2-4 24M 6F 20-40 Asian (Chinese) PS BF, SM
SFootBD [13, 10] 127 19,980 2 83M 44F 23.7 (18-32) Not Specified PS BF, SM
CAD WALK [11] 55 2,640 1 21M 34F 42.3 (18-70) White (Dutch) PS BF
UoM-Gait-69 [12] 69 14,394 3-4 32M 37F 36.4 (20-63) Not Specified PS, FS SO
UNB StepUP-P150 150 200,000 4-6 72M 78F 34.2 (19-91) White, Asian, Others PS, FS, SS, STS BF, SM, SC
@@
    b Downloadable resources only include data from 15 participants, of whom 13 wore two different shoe types, while the remaining 2 wore only one type.
    M: male, F: female, PS: preferred speed, FS: faster than preferred speed, SS: slower than preferred speed, STS: slow-to-stop, BF: barefoot, SO: personal shoes (one pair per participant), SM: personal shoes (multiple pairs per participant), SC: common shoes.
@@
None No spatial normalization.
Zero Padding Pad border of footstep with zeros to a specified tensor width and length.
Resize Apply spatial interpolation to resize the footstep to a specified tensor width and length.
Foot Rotation Rotate footstep according to the direction of its first principal component axis (the sole’s longest dimension).
@@
Registration Linearly transform (translate, rotate, scale) the footstep to align with a reference template (e.g., the MUN104 healthy barefoot templates [27]).
@@
None No temporal normalization.
Zero Padding Pad the 3D tensor with zeros at the end of the footstep to a specified number of frames.
Interpolation Interpolate the 3D tensor to a specified number of frames (e.g., 101 frames, where each frame represents 1% of the stance).
@@
None No amplitude normalization.
Body Mass Linearly rescale amplitudes by the participant’s measured body mass, so that amplitudes are relative to body weight.
Mean Pressure Linearly rescale amplitudes by the average total pressure of the footstep.
Min-Max Linearly rescale amplitudes to a maximum value of 1 and minimum of 0.
@@
ParticipantID 001-150 Unique participant identifier.
@@
Gender Man, Woman, Non-Binary, Not Specified Gender identity.
@@
[b] Field Possible Values/Format Description ParticipantID 001-150 Unique participant identifier. Footwear BF, ST, P1, P2 Footwear condition for trial. Speed W1, W2, W3, W4 Walking speed condition for trial. FootstepID* 0-$N_{steps}$ Footstep’s index in 90-second trial. PassID* 0-$N_{passes}$ The pass within which the footstep occurred (i.e., incremented each time the participant steps off the tiles to turn around). StartFrame* 0-$N_{frames}$ Frame in 90-second recording where footstep’s first pressure contact occurred. EndFrame* 0-$N_{frames}$ Frame in 90-second recording where footstep’s pressure contact ended. Ymin, Ymax* 0-719 Footstep bounding box limits along tile grid $y$-axis (parallel to walking direction, 3.6 m length). Xmin, Xmax* 0-239 Footstep bounding box limits along tile grid $x$-axis (perpendicular to walking direction, 1.2 m width). Orientation 0, 1 Footstep’s orientation on the tile grid: 1 if walking toward Tiles 11 and 12, or 0 if walking toward Tiles 1 and 2. Side Left, Right Whether the footstep corresponds to the right or left foot. Standing 0, 1 1 the footstep corresponds to standing behaviour during the Slow-to-Stop (W2) trials, 0 otherwise. Incomplete 0, 1 1 if the footstep was not captured fully by the sensors (e.g., partially off of the tile-grid or cut-off at beginning or end of recording), 0 otherwise. Rscore 0-89 Footstep’s R-score, approximating the number of standard deviations from the trial mean. Outlier 0, 1 1 if the footstep’s R-Score exceeds the recommended threshold of 2, or 0 otherwise. Exclude 0, 1 The combination of the ‘Standing’, ‘Incomplete’ and ‘Outlier’ columns for easy exclusion of these footsteps if desired. RotationAngle -90°- 90° Footstep’s rotation angle in degrees with respect to the tile grid’s long axis ($y$-axis). FootLength 1-75 Length of the sole’s contact area in pixels, measured along the footstep’s first principal component axis (longest foot dimension). FootWidth 1-40 Width of the sole’s contact area in pixels, measured along the footstep’s second principal component axis (perpendicular to longest foot dimension). MeanPressure Floating Point Average of the footstep’s total pressure (sum of all pressure values in kPa at each time point) over the duration of the stance.